# SciCoQA: Quality Assurance for Scientific Paper–Code Alignment

### 26 Mar 2026 v2

@@ 8 Ethical Considerations
We constructed the synthetic portion of SciCoQA using repositories with permissive licenses (MIT, Apache 2.0, BSD, CC-BY) to ensure respectful redistribution of code. For real-world discrepancies derived from GitHub issues and reproducibility reports, we leverage the publicly available data to create our ground truth data by contextualizing and rephrasing the original issue. None of our data contains any personal information; however, since we work with scientific publications, these are closely associated with the authors of the respective papers. We acknowledge that highlighting discrepancies in a specific author’s work may be perceived negatively. We emphasize that these discrepancies are treated as scientific artifacts for improving community reproducibility standards, rather than criticisms of individual researchers. Discrepancies frequently arise from benign causes, including concurrent code updates, simplifications for readability, or distinct authorship of paper and code, and their presence does not imply negligence or misconduct.
@@ Acknowledgments
We thank Jan Buchmann, Bhavyajeet Singh and Vatsal Venkatkrishna for their helpful feedback throughout the paper-writing process.
This work has been funded by the LOEWE Distinguished Chair “Ubiquitous Knowledge Processing”, LOEWE initiative, Hesse, Germany (Grant Number: LOEWE/4a//519/05/00.002(0002)/81) and by the German Federal Ministry of Research, Technology and Space and the Hessian Ministry of Higher Education, Research, Science and the Arts within their joint support of the National Research Center for Applied Cybersecurity ATHENE.
@@ A.2 Issue Processing
For the initial GitHub issue classification with Qwen3 4B Thinking, we used a low temperature of $0.2$ to minimize generation variance and promote strict adherence to the classification schema. For the verification of discrepancies from GitHub issues with GPT-5 we checked whether an issue description contained screenshots of the paper or code. If so, we replaced these with their text equivalent.
@@ A.3 Versioning
To ensure that the paper and code still contain the discrepancy, we provide links to versioned publications for arXiv, using the version at the time of the reporting of the discrepancy, i.e., publication of the reproducibility report or creation of the GitHub issue. Similarly, for the codebases, we use the commit history to obtain the version of the code at the time of the reporting of the discrepancy. Our dataset contains the versioned links to both the paper (for arxiv) and codebase. For the synthetic data, we take a version of the codebase at 31-10-2025.
A.4 Paper Processing
We provide as input the paper text in markdown. We use Mistrals OCR (Mistral AI Team, 2025) to convert the PDF to markdown. We exclude figures, although the captions are included. We further remove the references section from all papers.
@@ A.4 Paper Processing
GPT OSS 120B OpenAI (2025b) openai/gpt-oss-120b 131k Jun 2024 Aug 2025
GPT OSS 20B OpenAI (2025b) openai/gpt-oss-20b 131k Jun 2024 Aug 2025
GPT-5 OpenAI (2025a) gpt-5-2025-08-07 272k Sep 2024 Aug 2025
GPT-5 Codex OpenAI (2025a) gpt-5-codex 272k Sep 2024 Aug 2025
GPT-5 Mini OpenAI (2025a) gpt-5-mini-2025-08-07 272k May 2024 Aug 2025
GPT-5 Nano OpenAI (2025a) gpt-5-nano-2025-08-07 272k May 2024 Aug 2025
Devstral 24B Small Mistral AI and All Hands AI (2025) mistralai/Devstral-Small-2507 131k Mar 2025 Jul 2025
@@ A.5 Code Processing
To provide the code in the prompt, we obtain all files from the GitHub repository at a target date using the commit history, which we set to the date when the original GitHub issue was raised or the reproducibility paper was published. For the synthetic data, we set the cutoff at 31-10-2025. Having a versioned and static code repository is crucial to ensure reproducibility of our experiments and to make sure paper-code discrepancies have not been fixed in our dataset.
We filter the files to include only files that contain relevant content, mainly excluding dataset files. Specifically, we exclude files that are larger than 1MB and include files with the following extensions: .c, .cc, .cpp, .cu, .h, .hpp, .java, .jl, .m, .matlab, Makefile, .md, .pl, .ps1, .py, .r, .sh, config.txt, .rs, readme.txt, requirements_dev.txt, requirements-dev.txt, requirements.dev.txt, requirements.txt, .scala, .yaml, .yml. We obtain this list by manually inspecting the file extensions in the repositories of our dataset. We further process all jupyter notebook files to Python files by removing their output. For the prompt, we further include a list of all files in the repository to the model.
@@ A.7 Model and Decoding Configuration
We deploy the open-weight models with FP16, except for the GPT-OSS models, which were only published in MXFP4. We utilize the Ollama and vLLM libraries (Kwon et al., 2023) on 1-4 A100 80GB GPUs, depending on the model’s VRAM requirements; the commercial models are inferred via their respective APIs. We generally set the temperature to $1.0$, and for the GPT-5, GPT-OSS, and Gemini models, set a high reasoning effort/budget to maximize performance. For the Gemini models, we set a high reasoning budget of $24k$ tokens; for the GPT-5 models and GPT-OSS, we set a “high” reasoning budget. Further details about each model can be obtained from the respective model card linked in Table 2.
@@ B.1 Gemini and GPT Ground Truth Comparison
For the real data we rephrased the ground truth discrepancy descriptions with Gemini 3.1 Pro and GPT-5 (see §3.1). We analyze the similarity between these two descriptions in Fig. 7. We find the GPT descriptions to be slightly longer ($181\pm 40$ vs $135\pm 34$ tokens). We further find moderate lexical overlap between the two ground truths (average TF-IDF cosine similarity is $0.52$), but high semantic similarity (average $0.83$ computed using GemmaEmbedding), indicating that the models use different wording to describe the same discrepancy. Finally, we analyze how often the GPT-OSS 20B judge matches a prediction against none, either or both ground truths. We find that in $97.7\%$, the judge comes to the same verdict, given the two different discrepancy descriptions, yielding a Cohen’s Kappa of $\kappa=0.83$.
@@ B.2 Synthetic Code Analysis
We conduct a quantitative analysis of the synthetic data to validate that our generation pipeline produces minimal changes, leaving the main code intact. Fig. 6 presents the analysis of the generated code diffs.
@@
We find the generated discrepancies in the code to be highly targeted. The vast majority affect only a single file ($1.08\pm 0.29$ on average), with a maximum of three files. Similarly, the number of distinct generated diffs is low, with a mean of $1.24\pm 0.61$ snippets per discrepancy. This confirms that the generated errors are specific deviations in logic rather than large-scale refactors. Similarly, the number of line edits (i.e., added or removed lines) further supports the subtlety of the dataset. The distribution of line counts is heavily skewed towards small edits: On average, $2.41\pm 3.25$ lines are added and $3.13\pm 4.47$ lines are removed. The slightly higher count for removed lines suggests the discrepancies often simplify logic (e.g., removing a normalization step) rather than adding complex boilerplate.
@@
To quantify the similarity between the original and modified code, we calculate the Jaccard similarity. We measure the similarity between the entire generated code (which often contains a few more lines that are not changed), and only code lines that were modified. Comparing the entire code block generated by the model against the original yields a high mean similarity of $0.74\pm 0.20$. This indicates that the surrounding context and structure remain largely identical. Comparing only the specific lines that differ (added vs. removed) yields a mean similarity of $0.50\pm 0.29$. This moderate overlap suggests the changes are often modifications of existing statements (e.g., changing an operator or variable) rather than complete replacements. Overall, the data confirms that the synthetic discrepancies are precise and minimal, strictly adhering to the “small” and “conceptually meaningful” constraints of the synthetic data generation prompt (§F.5).
@@ C.1 Error Analysis
Table 3 provides the detailed breakdown of our human analysis from §5.2 on the predictions of GPT-5, Gemini 2.5 Pro and GPT-OSS 20B.
@@ C.2 Performance by Programming Language
In the synthetic dataset, the known injection points allow us to analyze detection performance across different programming languages. Since the dataset is skewed towards Python, drawing conclusions for specific low-resource languages is difficult due to the small sample size. Therefore, we group all non-Python instances into a unified subset of $46$ samples for analysis (“Not Python”). The resulting recall scores are presented in Table 4
Python vs. Non-Python Gap
We observe distinct behaviors across models regarding language robustness. Gemini 2.5 Pro and Flash models exhibit the largest performance drops on non-Python languages, with gaps of $10.2\%$ and $11.6\%$, respectively. However, other models are more stable, for example, GPT-5 or GPT-OSS 20B perform on par or better on the non-python discrepancies.
Language-Specific Performance
We highlight MATLAB as a positive outlier, where top models like GPT-5 and GPT-5 Mini achieve $90.0\%$ recall (significantly higher than their Python performance). We hypothesize that MATLAB’s mathematical syntax aligns more closely with the equations presented in scientific papers, facilitating easier visual and logical alignment for the models. Meanwhile, performance on systems-level languages like C/C++ and CUDA remains volatile and generally lower, though sample sizes limit definitive conclusions.
@@
Code File Prompt Position # Samples
$\leq 32k$ $253$ ($39.8\%$)
$\leq 64k$ $463$ ($72.9\%$)
$\leq 131k$ $568$ ($89.4\%$)
$\leq 262k$ $631$ ($99.4\%$)
$\leq 524k$ $635$ ($100.0\%$)
C.3 Truncation Analysis
In §5 we analyzed the model performance by input length, finding that longer inputs degrade performance. We extend this analysis by investigating the performance by the position of the last relevant code file in the prompt. The relevant code files are obtained in the validation/rephrasing step for the real data, and during discrepancy generation for the synthetic data. Table 5 shows the distribution of the position of the last relevant code file in the prompt. We find that $89.4\%$ of samples in our data contain all relevant code files within a context window of $131k$ tokens, and $99.4\%$ within a $262k$ context window.
We further analyze the recall by the position of the last relevant code file in the prompt and bin the samples and report results in Table 8. We observe substantial recall degradation even for models that experience virtually zero truncation. GPT-5 drops from $65.8\%$ to $50.7$ and Gemini 3.1 Pro from $55.0\%$ to $31.3\%$ as the relevant file appears later in the prompt.
Both the code file position distribution as well as the performance degradation demonstrate that the dominant factor is long-context reasoning, not truncation, consistent with our analysis in Fig. 4 and prior findings on long-context degradation (e.g. Liu et al. (2024b); Hsieh et al. (2024); Levy et al. (2024)).
C.4 Real vs Synthetic Recall
To assess the validity of our synthetic data generation pipeline, we analyze the relationship between model performance on real-world discrepancies versus synthetic discrepancies. Fig. 9 visualizes the recall and relationship between the subsets. We observe a high positive Pearson correlation of $r=0.94$. When excluding the GPT-5 family (i.e. GPT-5, GPT-5 Mini, GPT-5 Nano, and GPT-5 Codex) from the correlation analysis, the Pearson correlation rises to $r=0.98$. This indicates that the relative ranking of models remains consistent regardless of the data origin; models that demonstrate strong capabilities on our synthetic injections are reliably better at identifying real-world discrepancies. While this trend might be influenced by the general capabilities of the models the strong correlation confirms that the synthetic data does not introduce noise or bias that alters the model rankings. Consequently, the synthetic subset serves as a reliable proxy for evaluating performance, justifying its use for scaling the benchmark to scientific domains where real-world examples are scarce.
@@ C.4 Real vs Synthetic Recall
C.5 Gemini vs GPT Ground Truth Recall
A potential concern with the LLM-rephrased ground truth discrepancy descriptions is that a model from the same family might have an advantage when evaluated against the ground truth (GT) that was generated by the same model or from the same family. To test this, we compare per-model recall when evaluating against only the GPT-generated versus Gemini-generated ground truth descriptions across the $80$ discrepancies for which both are available (Table 7). We find no evidence of systematic provider bias. In fact, the closed OpenAI models perform the same or better against the Gemini-generated ground truth. Gemini 3.1 Pro performs slightly better against its own generations ($\Delta=-1.3$), but the other Gemini models do not exhibit any bias towards these descriptions. Among third-party model families, which serve as unbiased controls since they are not affiliated with either GT provider, results are similarly balanced: DeepSeek and NVIDIA models show no consistent preference, while the Qwen models slightly favor GPT GT (mean $\Delta=+2.0$pp). Across all $22$ models, the overall mean $\Delta$ is +0.2+0.2pp with a maximum absolute difference of just 3.73.7pp, and $9$ models favor GPT GT versus $8$ favoring Gemini GT ($5$ tied). These small, inconsistent differences suggest that the choice of ground truth source does not introduce meaningful evaluation bias.
C.6 Pooled Ground Truth Results
In §5.2 we have created a pooled ground truth for $20$ NLP and CV papers by manually verifying the predicted discrepancies from GPT-5, Gemini 2.5 Pro and GPT-OSS 20B. Including the existing discrepancies from the real-world data in SciCoQA on these papers, this resulted in a total of $129$ verified paper-code discrepancies. Following the pooling methodology from information retrieval (Buckley and Voorhees, 2004), where relevance judgments are aggregated from multiple systems to approximate a more complete ground truth, we evaluate all models against the pooled set. In case multiple ground truths exist for the same discrepancy, e.g. when both GPT-5 and Gemini 2.5 Pro surfaced the same unannotated discrepancy, we consider it detected if the prediction matches at least one description. Table 6 reports the results.
Model TP FP P R F1
GPT-5 $69$ $13$ $84.1$ $63.6$ $72.4$
GPT-5 Mini $70$ $29$ $70.7$ $62.8$ $66.5$
GPT-OSS 20B $71$ $32$ $68.9$ $63.6$ $66.1$
Gemini 2.5 Pro $54$ 99 $85.7$ $48.8$ $62.2$
Gemini 3.1 Pro $54$ $17$ $76.1$ $50.4$ $60.6$
GPT-OSS 120B $55$ $35$ $61.1$ $52.7$ $56.6$
Gemini 2.5 Flash $43$ $25$ $63.2$ $41.1$ $49.8$
GPT-5 Codex $34$ 99 $79.1$ $34.9$ $48.4$
Gemini 2.5 Flash Lite $28$ 2121 $57.1$ $30.2$ $39.5$
Nemotron 49B v1.5 $32$ $25$ $56.1$ $30.2$ $39.3$
GPT-5 Nano $31$ $39$ $44.3$ $34.1$ $38.5$
Devstral 24B Small $26$ $31$ $45.6$ $27.9$ $34.6$
Qwen3 30B Inst. $26$ $35$ $42.6$ $27.1$ $33.2$
Qwen3 30B Coder $27$ $42$ $39.1$ $27.9$ $32.6$
Qwen3 30B Think. 2020 $12$ $62.5$ $18.6$ $28.7$
Qwen3 4B Inst. 2121 $33$ $38.9$ $22.5$ $28.5$
DeepSeek R1 32B 2121 $43$ $32.8$ $22.5$ $26.7$
Magistral 24B Small $25$ $62$ $28.7$ $24.0$ $26.2$
Qwen3 4B Think. $16$ $10$ $61.5$ $14.7$ $23.8$
Nemotron Nano 9B v2 $23$ $75$ $23.5$ $17.8$ $20.3$
DeepSeek Coder 16B V2 22 $16$ $11.1$ $6.2$ $8.0$
DeepSeek R1 8B 22 $12$ $14.3$ $1.6$ $2.8$
Table 6: Precision (P), Recall (R) and F1 scores on 2020 NLP and CV papers with $129$ discrepancies. Ground truths have been pooled from predictions from GPT-5, Gemini 2.5 Pro and GPT-OSS 20B. Additionally, discrepancies from the real-world data of SciCoQA have been added.
GPT-5 achieves the highest F1 ($72.4\%$), balancing precision ($84.1\%$) and recall ($63.6\%$), followed by GPT-5 Mini and GPT-OSS 20B ($66.5\%$ and $66.1\%$). As established in the pooling literature, systems contributing to the pool have a structural advantage, so the evaluations of GPT-5, Gemini 2.5 Pro, and GPT-OSS 20B should be viewed as upper bounds (Buckley and Voorhees, 2004). Nonetheless, clear patterns emerge: a precision–recall tradeoff is evident across models, with Gemini 2.5 Pro reaching the highest precision ($85.7\%$) at the cost of recall ($48.8\%$). GPT-OSS 20B and 120B are the strongest open-weight model, while the remaining open-weight models lag considerably (next best: Nemotron 49B, F1 of $39.3\%$). Despite the pool bias favoring contributing systems, recall remains the primary bottleneck: even under favorable conditions, the best model detects at most $63.6\%$ of verified discrepancies.
@@ C.6 Pooled Ground Truth Results
Table 7: Model performance on the real-world SciCoQA data for the $80$ instances where both a GPT-5 and Gemini 3.1 Pro ground truth discrepancy description is available. The columns show the recall performance when assessed against both (Gemini+GPT), Gemini only or GPT only ground truth. A negative $\Delta$ shows preference for Gemini’s ground truth descriptions, while a positive $\Delta$ shows a preference for GPT’s.

### 22 Apr 2026 v3

@@ 8 Ethical Considerations
We constructed the synthetic portion of SciCoQA using repositories with permissive licenses (MIT, Apache 2.0, BSD, CC-BY) to ensure respectful redistribution of code. For real-world discrepancies derived from GitHub issues and reproducibility reports, we leverage the publicly available data to create our ground truth data by contextualizing and rephrasing the original issues. None of our data contains any personal information; however, since we work with scientific publications, they are closely associated with the authors of the respective papers. We acknowledge that highlighting discrepancies in a specific author’s work may be perceived negatively. We emphasize that these discrepancies are treated as scientific artifacts for improving community reproducibility standards, rather than criticisms of individual researchers. Discrepancies frequently arise from benign causes, including concurrent code updates, simplifications for readability, or distinct authorship of paper and code, and their presence does not imply negligence or misconduct.
@@ Acknowledgments
We thank Jan Buchmann, Bhavyajeet Singh, and Vatsal Venkatkrishna for their helpful comments throughout the paper-writing process, as well as the anonymous reviewers and area chair for their constructive feedback.
This work has been funded by the LOEWE Distinguished Chair “Ubiquitous Knowledge Processing,” LOEWE initiative, Hesse, Germany (Grant Number: LOEWE/4a//519/05/00.002(0002)/81) and by the German Federal Ministry of Research, Technology and Space and the Hessian Ministry of Higher Education, Research, Science and the Arts within their joint support of the National Research Center for Applied Cybersecurity ATHENE.
@@ A.2 Issue Processing
For the initial GitHub issue classification with Qwen3 4B Thinking, we used a low temperature of $0.2$ to minimize generation variance and promote strict adherence to the classification schema. For the verification of discrepancies from GitHub issues with GPT-5, we checked whether an issue description contained screenshots of the paper or code. If so, we replaced these with their text equivalents.
@@ A.3 Versioning
To ensure that the paper and code still contain the discrepancy, we provide links to versioned publications for arXiv, using the version at the time of the reporting of the discrepancy, i.e., publication of the reproducibility report or creation of the GitHub issue. Similarly, for the codebases, we use the commit history to obtain the version of the code at the time of the reporting of the discrepancy. Our dataset contains the versioned links to both the paper (for arXiv) and the codebase. For the synthetic data, we take a version of the codebase as of 31-10-2025.
@@
GPT-OSS 120B OpenAI (2025) openai/gpt-oss-120b 131k Jun 2024 Aug 2025
GPT-OSS 20B OpenAI (2025) openai/gpt-oss-20b 131k Jun 2024 Aug 2025
GPT-5 OpenAI (2025) gpt-5-2025-08-07 272k Sep 2024 Aug 2025
GPT-5 Codex OpenAI (2025) gpt-5-codex 272k Sep 2024 Aug 2025
GPT-5 Mini OpenAI (2025) gpt-5-mini-2025-08-07 272k May 2024 Aug 2025
GPT-5 Nano OpenAI (2025) gpt-5-nano-2025-08-07 272k May 2024 Aug 2025
Devstral 24B Small Mistral AI & All Hands AI (2025) mistralai/Devstral-Small-2507 131k Mar 2025 Jul 2025
@@
A.4 Paper Processing
We provide the paper text in markdown as input. We use Mistral’s OCR (Mistral AI Team, 2025) to convert the PDF to markdown. We exclude figures, although the captions are included. We further remove the references section from all papers.
@@ A.5 Code Processing
To provide the code in the prompt, we obtain all files from the GitHub repository at a target date using the commit history, which we set to the date when the original GitHub issue was raised or the reproducibility paper was published. For the synthetic data, we set the cutoff to 31-10-2025. Having a versioned and static code repository is crucial to ensure reproducibility of our experiments and to make sure paper-code discrepancies have not been fixed in our dataset.
We retain only files that contain relevant content, mainly excluding dataset files. Specifically, we exclude files that are larger than 1MB and include files with the following extensions: .c, .cc, .cpp, .cu, .h, .hpp, .java, .jl, .m, .matlab, Makefile, .md, .pl, .ps1, .py, .r, .sh, config.txt, .rs, readme.txt, requirements_dev.txt, requirements-dev.txt, requirements.dev.txt, requirements.txt, .scala, .yaml, .yml. We obtain this list by manually inspecting the file extensions in the repositories of our dataset. We further convert all Jupyter notebooks to Python files by removing their outputs. For the prompt, we further include a list of all files in the repository.
@@ A.7 Model and Decoding Configuration
We deploy the open-weight models with FP16, except for the GPT-OSS models, which were only published in MXFP4. We utilize the Ollama and vLLM libraries (Kwon et al., 2023) on 1-4 A100 80GB GPUs, depending on the model’s VRAM requirements; the commercial models are queried via their respective APIs. We generally set the temperature to $1.0$, and for the GPT-5, GPT-OSS, and Gemini models, set a high reasoning effort or budget to maximize performance. For the Gemini models, we set a high reasoning budget of $24k$ tokens; for the GPT-5 models and GPT-OSS, we set a “high” reasoning budget. Further details about each model can be obtained from the respective model card linked in Table 3.
@@ B.1 Gemini and GPT Ground Truth Comparison
For the real data, we rephrased the ground truth discrepancy descriptions with Gemini 3.1 Pro and GPT-5 (see §3.1). We analyze the similarity between these two descriptions in Fig. 6. We find the GPT descriptions to be slightly longer ($181\pm 40$ vs. $135\pm 34$ tokens). We further find moderate lexical overlap between the two ground truths (average TF-IDF cosine similarity is $0.52$), but high semantic similarity (average $0.83$ computed using EmbeddingGemma), indicating that the models use different wording to describe the same discrepancy. Finally, we analyze how often the GPT-OSS 20B judge matches a prediction against none, either, or both ground truths. We find that in $97.7\%$ of cases, the judge comes to the same verdict given the two different discrepancy descriptions, yielding a Cohen’s Kappa of $\kappa=0.83$.
@@ B.2 Synthetic Code Analysis
We conduct a quantitative analysis of the synthetic data to validate that our generation pipeline produces minimal changes, leaving the main code intact. Fig. 7 presents the analysis of the generated diffs.
@@
We find the generated discrepancies in the code to be highly targeted. The vast majority affect only a single file ($1.08\pm 0.29$ on average), with a maximum of three files. Similarly, the number of distinct generated diffs is low, with a mean of $1.24\pm 0.61$ snippets per discrepancy. This confirms that the generated errors are specific deviations in logic rather than large-scale refactors. The number of line edits (i.e., added or removed lines) further supports the subtlety of the dataset. The distribution of line counts is heavily skewed towards small edits: On average, $2.41\pm 3.25$ lines are added and $3.13\pm 4.47$ lines are removed. The slightly higher count for removed lines suggests the discrepancies often simplify logic (e.g., removing a normalization step) rather than adding complex boilerplate.
@@
To quantify the similarity between the original and modified code, we calculate the Jaccard similarity. We measure the similarity at two levels: over the entire generated code block (which often contains a few unchanged lines) and over only the code lines that were modified. Comparing the entire code block generated by the model against the original yields a high mean similarity of $0.74\pm 0.20$. This indicates that the surrounding context and structure remain largely identical. Comparing only the specific lines that differ (added vs. removed) yields a mean similarity of $0.50\pm 0.29$. This moderate overlap suggests the changes are often modifications of existing statements (e.g., changing an operator or variable) rather than complete replacements. Overall, the data confirms that the synthetic discrepancies are precise and minimal, strictly adhering to the “small” and “conceptually meaningful” constraints of the synthetic data generation prompt (§F.5).
@@ C.1 Error Analysis
Table 4 provides the detailed breakdown of our human analysis from §5.2 on the predictions of GPT-5, Gemini 2.5 Pro, and GPT-OSS 20B.
@@ C.2 Performance by Programming Language
In the synthetic dataset, the known injection points allow us to analyze detection performance across different programming languages. Since the dataset is skewed towards Python, drawing conclusions for specific low-resource languages is difficult due to the small sample size. Therefore, we group all non-Python instances into a unified subset of $46$ samples for analysis (“Not Python”). The resulting recall scores are presented in Table 6. Most models show only small, inconsistent differences between Python and non-Python recall (within $\pm 5$ pp). The notable exceptions are Devstral 24B Small, which collapses from $15.1\%$ on Python to $0\%$ on non-Python, and Gemini 2.5 Flash (+9.7+9.7 pp); conversely, GPT-5 Mini (5.9-5.9 pp), GPT-OSS 20B (5.4-5.4 pp), and Nemotron Nano 9B v2 (7.8-7.8 pp) actually perform slightly better on non-Python. Beyond these aggregate trends, two specific languages stand out. MATLAB is a positive outlier: GPT-5 and GPT-5 Mini both achieve $90.0\%$ recall, substantially above their Python performance. With $n=10$ this is tentative, but we hypothesize that MATLAB’s mathematical syntax aligns more closely with the equations presented in scientific papers, facilitating easier logical alignment for the models. Conversely, CUDA is a near-universal weak spot: $17$ of $22$ models score $0\%$ recall on CUDA ($n=4$), including GPT-5, Gemini 3.1 Pro, and GPT-OSS 120B. Small sample sizes limit definitive conclusions for both languages.
Code File Prompt Position # Samples
$\leq 32k$ $253$ ($39.8\%$)
$\leq 64k$ $463$ ($72.9\%$)
$\leq 131k$ $568$ ($89.4\%$)
$\leq 262k$ $631$ ($99.4\%$)
$\leq 524k$ $635$ ($100.0\%$)
C.3 Truncation Analysis
@@
In §5 we analyzed the model performance by input length, finding that longer inputs degrade performance. We extend this analysis by investigating the performance as a function of the position of the last relevant code file in the prompt. The relevant code files are obtained in the validation/rephrasing step for the real data, and during discrepancy generation for the synthetic data. Table 5 shows the distribution of the position of the last relevant code file in the prompt. We find that $89.4\%$ of samples in our data contain all relevant code files within a context window of $131k$ tokens, and $99.4\%$ within a $262k$ context window.
We further analyze recall by the position of the last relevant code file in the prompt, binning the samples and reporting results in Table 9. We observe substantial recall degradation even for models that experience virtually zero truncation. GPT-5 drops from $77.5\%$ to $50.7\%$ and Gemini 3.1 Pro from $64.4\%$ to $37.3\%$ as the relevant file appears later in the prompt.
Both the code file position distribution and the performance degradation demonstrate that the dominant factor is long-context reasoning, not truncation, consistent with our analysis in Fig. 4 and prior findings on long-context degradation (Liu et al., 2024b; Hsieh et al., 2024; Levy et al., 2024).
@@
C.4 Real vs. Synthetic Recall
@@
Table 7: Model performance on the real-world SciCoQA data for the $80$ instances where both a GPT-5 and Gemini 3.1 Pro ground truth discrepancy description is available. The columns show the recall performance when assessed against both (Gemini+GPT), Gemini-only, or GPT-only ground truth. A negative $\Delta$ shows a preference for Gemini’s ground truth descriptions, while a positive $\Delta$ shows a preference for GPT’s.
To assess the validity of our synthetic data generation pipeline, we analyze the relationship between model performance on real-world discrepancies versus synthetic discrepancies. Fig. 9 visualizes the recall and the relationship between the subsets. We observe a high positive Pearson correlation of $r=0.94$. When excluding the GPT-5 family (i.e., GPT-5, GPT-5 Mini, GPT-5 Nano, and GPT-5 Codex) from the correlation analysis, the Pearson correlation rises to $r=0.98$, indicating that the GPT-5 family sits slightly off the shared trend line that the other models follow closely. Apart from this caveat, the relative ranking of models remains consistent regardless of the data origin; models that demonstrate strong capabilities on our synthetic injections are reliably better at identifying real-world discrepancies. While this trend might be influenced by the general capabilities of the models, the strong correlation confirms that the synthetic data preserves relative model rankings. Consequently, absolute performance should be assessed on the real subset, while the synthetic subset serves as a reliable proxy for relative model comparison, justifying its use for scaling the benchmark to scientific domains where real-world examples are scarce.
C.5 Gemini vs. GPT Ground Truth Recall
A potential concern with the LLM-rephrased ground truth discrepancy descriptions is that a model from the same family might have an advantage when evaluated against the ground truth (GT) that was generated by the same model or from the same family. To test this, we compare per-model recall when evaluating against only the GPT-generated versus Gemini-generated ground truth descriptions across the $80$ discrepancies for which both are available. We report results in Table 7. We find no evidence of systematic provider bias. In fact, the closed OpenAI models perform the same or better against the Gemini-generated ground truth. Gemini 3.1 Pro performs slightly better against its own generations ($\Delta=-1.3$), but the other Gemini models do not exhibit any bias towards these descriptions. Among third-party model families, which serve as unbiased controls since they are not affiliated with either GT provider, results are similarly balanced: DeepSeek and NVIDIA models show no consistent preference, while the Qwen models slightly favor GPT GT (mean $\Delta=+2.0$pp). Across all $22$ models, the overall mean $\Delta$ is +0.2+0.2pp with a maximum absolute difference of just 3.73.7pp, and $9$ models favor GPT GT versus $8$ favoring Gemini GT ($5$ tied). These small, inconsistent differences suggest that the choice of ground truth source does not introduce meaningful evaluation bias.
C.6 Pooled Ground Truth Results
Model P R F1
GPT-5 $84.1$ $63.6$ $72.4$
GPT-5 Mini $70.7$ $62.8$ $66.5$
GPT-OSS 20B $68.9$ $63.6$ $66.1$
Gemini 2.5 Pro $85.7$ $48.8$ $62.2$
Gemini 3.1 Pro $76.1$ $50.4$ $60.6$
GPT-OSS 120B $61.1$ $52.7$ $56.6$
Gemini 2.5 Flash $63.2$ $41.1$ $49.8$
GPT-5 Codex $79.1$ $34.9$ $48.4$
Gemini 2.5 Flash Lite $57.1$ $30.2$ $39.5$
Nemotron Super 49B v1.5 $56.1$ $30.2$ $39.3$
GPT-5 Nano $44.3$ $34.1$ $38.5$
Devstral 24B Small $45.6$ $27.9$ $34.6$
Qwen3 30B Inst. $42.6$ $27.1$ $33.2$
Qwen3 30B Coder $39.1$ $27.9$ $32.6$
Qwen3 30B Think. $62.5$ $18.6$ $28.7$
Qwen3 4B Inst. $38.9$ $22.5$ $28.5$
DeepSeek R1 32B $32.8$ $22.5$ $26.7$
Magistral 24B Small $28.7$ $24.0$ $26.2$
Qwen3 4B Think. $61.5$ $14.7$ $23.8$
Nemotron Nano 9B v2 $23.5$ $17.8$ $20.3$
DeepSeek Coder 16B V2 $11.1$ $6.2$ $8.0$
DeepSeek R1 8B $14.3$ $1.6$ $2.8$
Table 8: Precision (P), Recall (R), and F1 scores on 2020 NLP and CV papers with $129$ discrepancies. Ground truths have been pooled from the predictions of GPT-5, Gemini 2.5 Pro, and GPT-OSS 20B. Additionally, discrepancies from the real-world data of SciCoQA have been added.
In §5.2 we created a pooled ground truth for $20$ NLP and CV papers by manually verifying the predicted discrepancies from GPT-5, Gemini 2.5 Pro, and GPT-OSS 20B. Together with the existing discrepancies from the real-world data in SciCoQA on these papers, this yields $129$ verified paper-code discrepancies. Following the pooling methodology from information retrieval (Buckley and Voorhees, 2004), where relevance judgments are aggregated from multiple systems to approximate a more complete ground truth, we evaluate all models against the pooled set. In case multiple ground truths exist for the same discrepancy, e.g., when both GPT-5 and Gemini 2.5 Pro surfaced the same unannotated discrepancy, we consider it detected if the prediction matches at least one description. Table 8 reports the results.
GPT-5 achieves the highest F1 ($72.4\%$), balancing precision ($84.1\%$) and recall ($63.6\%$) the best, followed by GPT-5 Mini and GPT-OSS 20B ($66.5\%$ and $66.1\%$). As established in the pooling literature, systems contributing to the pool have a structural advantage, so the evaluations of GPT-5, Gemini 2.5 Pro, and GPT-OSS 20B should be viewed as upper bounds (Buckley and Voorhees, 2004). Nonetheless, clear patterns emerge: a precision–recall tradeoff is evident across models, with Gemini 2.5 Pro reaching the highest precision ($85.7\%$) at the cost of recall ($48.8\%$). GPT-OSS 20B and 120B are the strongest open-weight models, while the remaining open-weight models lag behind considerably (next best: Nemotron 49B, F1 of $39.3\%$). Despite the pool bias favoring contributing systems, recall remains the primary bottleneck: even under favorable conditions, the best model detects at most $63.6\%$ of verified discrepancies.